OpenStudy (anonymous):

For the following equation, state the value of the discrminant and then describe the nature of solutions. -7x^2-10x+11=0 a. the equation has 2 real solutions. b. the equation has 1 real solution. c. the equation has 2 imaginary solutions.

OpenStudy (anonymous):

is it 102 as the discriminant and c is the answer?

OpenStudy (anonymous):

i think the discriminant is 408

OpenStudy (anonymous):

\[(-10)^2-4\times (-7) \times 11=100+308=408\]

OpenStudy (anonymous):

still get two real solutions though, since it is postive

OpenStudy (anonymous):

hmm...ok, thanks...can you tell me how to get the descriminant, i was trying to look at my book, and its harder to understand..so i was trying to google it..confused here.

OpenStudy (anonymous):

i see your solution...

OpenStudy (anonymous):

sure discriminant for \[ax^2+bc+c=0\] is \[b^2-4ac\]

OpenStudy (anonymous):

sorry i meant for \[ax^2+bx+c=0\]

OpenStudy (anonymous):

I saw that part on how to do it..got confused..then i was looking at the purple math, more confused i got..i think i just need a break lol...thanks for the help satellite73...will be back shortly.

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thanks bunches...:)